artlover03:

A construction crew is lengthening a road. The road started with a length of 52 miles, and the crew is adding 3 miles to the road each day. Let L represent the total length of the road (in miles), and let D represent the number of days the crew has worked. Write an equation relating L to D. Then use this equation to find the total length of the road after the crew has worked 31 days

artlover03:

@MrMudd183

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so first we need to identify the variables. As said in the question what are the variables? \(\color{#0cbb34}{\text{End of Quote}}\) (Asking you what the variables are)

artlover03:

L is length in road in miles and D represents the # of days the crew worked

justus:

that's correct

supie:

Yeah, that's correct. @justus u wanna take it from here?

justus:

now you need to find the numbers for each variable.

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Yeah, that's correct. @justus u wanna take it from here? \(\color{#0cbb34}{\text{End of Quote}}\) Sure :)

justus:

what was the initial length? and what was the number of days?

artlover03:

52 to begin with while adding 3 miles x day

justus:

so it would be L = 52 + 2D, now they are asking for 31 days so we would have to plug it in the equation then it would be L = 52 + 2(31) =

artlover03:

L=114 ?

justus:

that's right.

justus:

thank you for being an understanding and cooperating user.